OpenStudy (anonymous):

solve cos^2 x-5sin x +5=0

OpenStudy (anonymous):

step by step please

OpenStudy (anonymous):

convert cos^2 x into sin^2 x then take sin x =t solve the quad then substitue t =sin x AND get ur soln AM NOT GIVING OUT THE WHOLE SOLUTION

OpenStudy (anonymous):

\[1-\sin ^{2}x - 5 \sin x + 5\]

OpenStudy (anonymous):

u r going r8

OpenStudy (anonymous):

then, \[-\sin ^2 x -5 sinx +6\]

OpenStudy (anonymous):

then, \[-(sinx - 1)(\sin x+6)\]

OpenStudy (anonymous):

then you solve sin x =-1 and sin x=6 right?

OpenStudy (anonymous):

which would be 2pi and pi?

OpenStudy (anonymous):

right and sinx cant be 6 implies sinx to be 1

OpenStudy (anonymous):

im lost there

OpenStudy (anonymous):

why does it do that?

OpenStudy (anonymous):

coz sin graph is confined between -1 and 1

OpenStudy (anonymous):

okay, so if it is 6 i thought it was undefined...

OpenStudy (anonymous):

yess u did that good

OpenStudy (anonymous):

so then why would sin x =6 be 1?

OpenStudy (anonymous):

sin x could be 1 or 6...... we cant choose 6 as it solution coz sin has range of [-1,1] so the only soln left is sin x=1

OpenStudy (anonymous):

now u get it?

OpenStudy (anonymous):

so it would be Pi/2 +2npi

OpenStudy (anonymous):

but at the beginning of it you have -(sin x-1)(sin x +6) what do u do with the negative in the front?

OpenStudy (anonymous):

general soln is diff from this

OpenStudy (anonymous):

because sin x can't equal 6 so that would be thrown out

OpenStudy (anonymous):

idk what you mean

OpenStudy (anonymous):

no prob of negative sign coz of zero on other side

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

when we write soln in term of n... then its called the general soln and that is diff from urs

OpenStudy (anonymous):

so??? it would be what? just +npi? or just pi/2

OpenStudy (anonymous):

x = nπ + (-1)ⁿ.pi/2

OpenStudy (anonymous):

why would it be a negative pi/2

OpenStudy (anonymous):

i dont get that

OpenStudy (anonymous):

what you are saying

OpenStudy (anonymous):

it looks to me what you just sent me it is -1pi/2

OpenStudy (anonymous):

or is it just pi/2+npi

OpenStudy (anonymous):

its n (pi) + (-1)^n pi/2

OpenStudy (anonymous):

and n is a whole no.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so yes or no your answer is pi/2+npi

OpenStudy (anonymous):

ques clear?

OpenStudy (anonymous):

my answer is already stated above

OpenStudy (anonymous):

so what do u say?